please help me find this. A rectangle has an area of 12 square inches and a perimeter of 16 square inches. Find the length and width of the rectangle. Show your work.
Answer by stanbon(75887)

please help me find this. A rectangle has an area of 12 square inches and a perimeter of 16 square inches. Find the length and width of the rectangle. Show your work.
----------------------------
Perimeter = 2(length + width)
Area = length*width
---------------------------
Equations:
L + W = 8
LW = 12
-----
Solve by substitution:
L = 8-W
---
Substitute for "L" in the 2nd equation and solve for "W":
(8-W)W = 12
8W-W^2 = 12
---
Rearrange:
W^2 - 8W +12 = 0
---
Factor:
(W-6)(W-2) = 0
---
If W = 6, L = 2
If W = 2, L = 6
===================
Cheers,
Stan H.